### v1

# Quantum Algorithm of the GLMY Homology on Digraphs

Abstract
Quantum algorithms for topological data analysis provide significant advantage over the best classical algorithm. Different from the previous simplical complex on points cloud, the GLMY homology introduced by Alexander Grigor’yan, Yong Lin, Yuri Muranov and Shing-Tung Yau [1], is defined on digraph and is a arising realm in Topological Data Analysis (TDA), which attracts more and more attention recently. We propose a quantum algorithm for the GLMY homology with significant advantage over the best classical algorithm. We design a universal encoding protocol for the quantum states and boundary operators of GLMY homology on digraphs. And a property of the GLMY homology is proved for the theoretical guarantee of the quantum algorithm. The quantum algorithm for GLMY homology gives a quadratic speedup in general cases, and it gives an exponential quantum advantage in the case of the input data is given as a specification of paths.
I Introduction
Topological Data Analysis (TDA) is a mathematical field focus on the ‘shapes’ of the data. By applying the topological tools, for example the simplicial homology theory and the knot theory etc., it is able to analyze complex high-dimensional data and identify key features of the data which may be difficult to be extracted with traditional methods. Quantum algorithm for TDA is first proposed by Lloyd, Garnerone, and Zanardi (LGZ) [2], and it gives significant advantage over the best classical algorithm. The foundations of the LGZ algorithm are linear-algebraic quantum machine learning (QML) algorithms [3, 4, 5, 6, 7, 8, 9, 10], which are one of the most rapidly growing fields in the research of quantum computing. Recent studies on quantum algorithms for TDA have achieved notable theoretical progress [11, 12, 13, 14, 15, 16, 17, 18]. The complexity of the LGZ algorithm is analyzed in [17] and the conclusion is that computing Betti numbers is NP-hard and it achieves quadratic speedup for inputs as vertices and edges. It can achieve exponential advantage only when inputs are given as a simplex list. The LGZ algorithms are also experimentally implemented on optical platform [19] and noisy quantum simulator [20].
The path homology, which is also known as GLMY homology, is a arising realm in TDA that attracting more and more attention recently. It was first introduced and named by Alexander Grigor’yan, Yong Lin, Yuri Muranov and Shing-Tung Yau in the year 2012 within the paper [1]. This is a theory describing the topological properties of a given digraph. Mathematicians has developed in recent years, the homotopy theory based on the GLMY homolgy in [21], the GLMY cohomology theory in [22], the persistent GLMY homolgy in [23] and the analytic and Reidemeister torsions of digraphs in [24]. Moreover the GLMY homology theory has been generalized to describe the topological properties of multigraphs and Quivers in [25], weighted digraphs in [26] and weighted hypergraphs in [27]. The GLMY homology also showed potential power in the applied mathematics, for example the applications in deep feedback networks in [28], temporal networks in [29], molecular and materials sciences in [30]. However in the studies and applications of this theory, computation of the GLMY homology group was always a shadow above the Mathematicians. Some attempts for accelerating the computation has made for example [31] and [32]. However the best known algorithm for computing $k$-th GLMY homology was the one given in [23] which has time complexity $\mathcal{O}(n^{3+3k})$ providing $n$ the number of vertices in the digraph.
Here we propose a quantum algorithm for the GLMY homology on digraphs. The framework is based on the LGZ algorithm, which incorporates three components: the preparation of input states, Hamiltonian simulation of the Laplacian operator and phase estimate of the eigenvalues. We design a quantum state encoding protocol for digraphs based on paths in the GLMY homology. And we design a encoding protocol of the boundary map for quantum circuits. Our encoding protocol is universal for the paths on digraphs. Furthermore, a property of the GLMY homology is proved for the theoretical guarantee of the quantum algorithm thus the norm matrix used in classical algorithm can be ignored. Given the input data as a specification of paths, the amplitude encoding method can be used to efficiently prepare the input state. The quantum algorithm for GLMY homology gives a quadratic speedup in general cases, and it gives an exponential quantum advantage in the special case of the input data is given as a specification of paths that can be sampled efficiently.
II GLMY Homology: Definitions and Computations
A digraph $G=(V,E)$ is a tuple where $V=\{1,\cdots,n\}$ is the set of vertices, and $E$ is a finite set of ordered pairs $(v_{i}\to v_{j})$ with $v_{i},v_{j}\in V$ and $v_{i}\neq v_{j}$, called the set of arrows or the set of directed edges. For simplicity, we will always assume that there is no loops in the digraph in the following discussion, i.e. there exists no sequence of arrows $v_{0}\to v_{1},v_{1}\to v_{2},\cdots,v_{k-1}\to v_{k}$ where $k$ is a positive integer such that $v_{0}=v_{k}$.
Figure 1: The illustration of digraph of GLMY homology and the boundary operator $\partial$ for GLMY homology. (a) Two digraphs with $n=4$ and $n=6$ and their Betti numbers. (b) The illustration of the boundary operator $\partial$ acting on paths.
II.1 Definitions
An elementary $k$-path is an ordered chain of vertices denoted as $e_{v_{0}\cdots v_{k}}:=v_{0}\cdots v_{k}$, where the number $k$ is called the length of the path. An elementary $k$-path $v_{0}\cdots v_{k}$ is called regular if and only if $v_{i}\neq v_{i+1}$ for all $i\in\{0,\cdots,k-1\}$ and irregular otherwise. The space spanned by all the regular elementary paths of length $k$ in one digraph is denoted as $\Lambda_{k}$. If one such regular elementary $k$-path satisfying the condition $(v_{i}\to v_{i+1})\in E$, we call it allowed. The space spanned by all allowed elementary $k$-paths in $E$ is denoted as $\mathcal{A}_{k}$. The total boundary map $\partial_{k}:\Lambda_{k}\to\Lambda_{k-1}$, which is given on each elementary $k$-paths and extended linearly, is defined by
$\partial_{k}(v_{0}\cdots v_{k})=\sum_{i=0}^{k}(-1)^{i}v_{0}\cdots\hat{v}_{i}\cdots v_{k},$ (1)
where $\hat{v}_{i}$ means the vertex $v_{i}$ is removed from the path. We call $\partial_{k}$ the $k$-th boundary map. The illustrations of the paths on a digraph and the boundary map are plotted in Fig. 1.
A standard approach for constructing the GLMY Homology is by taking the chain complex as the vector space $\Omega_{k}=\{x\in\mathcal{A}_{k}:\partial_{k}x\in\mathcal{A}_{k-1}\}$ and restricting the boundary maps $\partial_{k}$ to the space $\Omega_{k}$, denoted as $\partial^{\Omega}_{k}$. It is clear the spaces $(\Omega_{k},\partial^{\Omega}_{k})$ constitute a long chain with $\partial^{\Omega}_{k-1}\circ\partial^{\Omega}_{k}=0$,
$\Omega_{k}\stackrel{{\scriptstyle\partial^{\Omega}_{k}}}{{\longrightarrow}}\Omega_{k-1}\stackrel{{\scriptstyle\partial^{\Omega}_{k-1}}}{{\longrightarrow}}\dots\stackrel{{\scriptstyle\partial^{\Omega}_{1}}}{{\longrightarrow}}\Omega_{0}\stackrel{{\scriptstyle\partial^{\Omega}_{0}}}{{\longrightarrow}}0.$ (2)
And the $k$-th GLMY homology group $H_{k}(G)$ is defined as $\text{Ker}\partial^{\Omega}_{k}/\text{Im}\partial^{\Omega}_{k+1}$. Practically, the computation of $\Omega_{k}$ and $\partial^{\Omega}_{k}$ could be difficult in classical computations since it is based on solving linear equations which suffers with high complexities.
There is an alternative definition of GLMY homology which is known as the embedded homology group and proved to be isomorphic to the GLMY homology [33]. Let
$\Gamma_{k}:=\mathcal{A}_{k}+\partial_{k+1}\mathcal{A}_{k+1}.$ (3)
The inclusion $e_{k}:\Gamma_{k}\subset\Lambda_{k}$ is an embedding, therefore we take the boundary map $\partial^{\Gamma}_{k}$ as the restriction of $\partial_{k}$ on the space $\Gamma_{k}$. Clearly this boundary map satisfies the equation
$e_{k-1}\circ\partial_{k}^{\Gamma}=\partial_{k}\circ e_{k}.$ (4)
There is a chain
$\Gamma_{k}\stackrel{{\scriptstyle\partial^{\Gamma}_{k}}}{{\longrightarrow}}\Gamma_{k-1}\stackrel{{\scriptstyle\partial^{\Gamma}_{k-1}}}{{\longrightarrow}}\dots\stackrel{{\scriptstyle\partial^{\Gamma}_{1}}}{{\longrightarrow}}\Gamma_{0}\stackrel{{\scriptstyle\partial^{\Gamma}_{0}}}{{\longrightarrow}}0.$ (5)
The $k$-th embedded homology $\mathbf{H}_{k}$ is defined as the quotient group,
$\text{Ker }\partial^{\Gamma}_{k}/\text{Im }\partial^{\Gamma}_{k+1}.$ (6)
The dimensions of the homology groups $\mathbf{H}_{k}$ are important topological invariants, which is named as the $k$-th Betti number $\beta_{k}$ of the GLMY homology of $G$. Betti numbers are also the invariants widely used in the topological data analysis.
II.2 The Hodge Laplacian Operator
There exists a naturally defined inner product $\langle-,$-$\rangle$ on the space $\Lambda_{k}$, for any two elementary $k$-paths $v_{0}\cdots v_{k}$ and $u_{0}\cdots u_{k}$, $\langle v_{0}\cdots v_{k},u_{0}\cdots u_{k}\rangle$ equals to 1 if and only if $v_{0}=u_{0},v_{1}=u_{1},\cdots,v_{k}=u_{k}$ and 0 otherwise. The subspace $\Gamma_{k}$ inherits this inner strucutre from $\Lambda_{k}$ which is denoted as $\langle-,-\rangle_{\Gamma}$. The dual map of $\partial_{k}$ resp. $\partial^{\Gamma}_{k}$ with respect to $\langle-,$-$\rangle$ resp. $\langle-,-\rangle_{\Gamma}$ denoted by $\partial_{k}^{*}:\Lambda_{k-1}\to\Lambda_{k}$ resp. $(\partial^{\Gamma}_{k})^{*}:\Gamma_{k-1}\to\Gamma_{k}$, is given by the following relations,
$\langle\partial_{k}x,y\rangle$ $=$ $\ \langle x,(\partial_{k})^{*}y\rangle,$ (7)
$\langle\partial^{\Gamma}_{k}x^{\prime},y^{\prime}\rangle_{\Gamma}$ $=$ $\ \langle x^{\prime},(\partial^{\Gamma}_{k})^{*}y^{\prime}\rangle_{\Gamma},$
where $x\in\Lambda_{k},\ x^{\prime}\in\Gamma_{k},y\in\Lambda_{k-1}$ and $y^{\prime}\in\Gamma_{k-1}$ are arbitrary $k$-paths and (k1)(k-1)-paths. Their relations are given in the following Theorem, and the proof is left in Appendix A.
Theorem 1.
The dual maps $\partial_{k}^{*}$ and $(\partial^{\Gamma}_{k})^{*}$ satisfies the relation that $e_{k}\circ(\partial_{k}^{\Gamma})^{*}=\partial_{k}^{*}\circ e_{k-1}$, i.e. the following diagram commutes,
${\Gamma_{k}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{e_{k}}$${\Gamma_{k-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\partial_{k}^{\Gamma})^{*}}$$\scriptstyle{e_{k-1}}$${\Lambda_{k}}$${\Lambda_{k-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces.}$$\scriptstyle{\partial_{k}^{*}}$ (8)
The $k$-th total Hodge-Laplacian operator $\Delta_{k}:\Lambda_{k}\to\Lambda_{k}$ with respect to the inner product $\langle-,$-$\rangle$ and the $k$-th Hodge-Laplacian operator $\Delta^{\Gamma}_{k}:\Gamma_{k}\to\Gamma_{k}$ with respect to the inner product $\langle-,-\rangle_{\Gamma}$ is given as,
$\Delta_{k}=(\partial_{k})^{*}\partial_{k}+\partial_{k+1}(\partial_{k+1})^{*},$ (9)
$\Delta^{\Gamma}_{k}=(\partial^{\Gamma}_{k})^{*}\partial^{\Gamma}_{k}+\partial^{\Gamma}_{k+1}(\partial^{\Gamma}_{k+1})^{*}.$ (10)
It is well known [24] that the $k$-th Betti number of the GLMY homology group of $G$ is computed as
$\beta_{k}=\text{dim}\ \text{Ker}\partial^{\Gamma}_{k}/\text{Im}\partial^{\Gamma}_{k+1}=\text{dim}\ \text{Ker}\Delta^{\Gamma}_{k}.$ (11)
In Appendix D, we provide two examples of the classical computation of GLMY Homology. Moreover generally the space $\Gamma_{k}$ could be factorized as the direct sum below called the Hodge decomposition.
$\Gamma_{k}=\text{Ker}\Delta^{\Gamma}_{k}\oplus\partial^{\Gamma}_{k+1}\Gamma_{k+1}\oplus(\partial^{\Gamma}_{k-1})^{\ast}\Gamma_{k-1}.$ (12)
The following theorem is crucial for the quantum algorithm. And the proof of the Theorem is left in Appendix B.
Theorem 2.
The $k$-th Hodge-Laplacian operator $\Delta^{\Gamma}_{k}$ has the same nullspace as the operator $p_{k}\circ(\Delta_{k}|_{\Gamma_{k}})$ where $p_{k}$ is the projection from $\Lambda_{k}$ to $\Gamma_{k}$.
III Quantum Algorithm of GLMY Homology
In this section we will give the quantum algorithm of computing the Betti numbers of GLMY homology of a digraph $G$. In the quantum algorithm, the Hilbert space is identified with the vector space endowed with inner product discussed in Sec. II.2. In particular, the Hodge-Laplacian operator $\Delta_{k}$ or $\Delta_{k}^{\Gamma}$ are Hermitian operators defined on subspace $\Lambda_{k}$ or $\Gamma_{k}$, and the $k$-th Betti number is related with the degeneracy of $\Delta_{k}^{\Gamma}$ with eigen value being 0.
The digraph $G$ we consider has $n$ vertices labeled by $\{1,\cdots,n\}$, and $d$ ($d\leq n-1$) is the maximal path length in $G$. In this case we only need to consider chain groups from $\Gamma_{0}$ up to $\Gamma_{d}$. Since $\partial^{\Gamma}_{d}\Gamma_{d}=\partial^{\Gamma}_{d}\mathcal{A}_{d}$, we may replace the highest term $\Gamma_{d}$ with $\mathcal{A}_{d}$ in our algorithm.
Following the framework of the LGZ algorithm, a standard quantum algorithm of computing the $k$-th Betti number of GLMY homology consists of three main steps. The first step is to prepare a mixed state $u_{k}$ of the $k$-path simplices set in $\Gamma_{k}$
$u_{k}=\frac{1}{\dim\Gamma_{k}}\sum_{p_{k}\in\Gamma_{k}}|p_{k}\rangle\langle p_{k}|,$ (13)
and the vector $|p_{k}\rangle$ is the quantum state of the elementary $k$-path $p_{k}\equiv v_{0}...v_{k}$.
The second step is to encode the combinatorial Laplacian operator of Eq. (9),
$\Delta=\mathop{\bigoplus}\limits_{k=0}^{n-1}\Delta_{k}=\begin{pmatrix}\Delta_{0}&0&\cdots&0\\ 0&\Delta_{1}&&0\\ \vdots&&\ddots&\vdots\\ 0&\cdots&\cdots&\Delta_{{n-1}}\end{pmatrix}$. (14)
In the third step, we implement the unitary matrix $e^{i\Delta}$ to perform the quantum phase estimation algorithm with the input state $u_{k}$. Next we repeat the phase estimation algorithm multiple times. The kernel of the $\Delta^{\Gamma}$ is estimated based on the relative frequency of the measured eigenvalues being zero.
According to Eq. (11), to estimate the Betti numbers, we need to implement the transformation $e^{i\Delta^{\Gamma}}$ instead of $e^{i\Delta}$. However, by Theorem 2 and the following Theorem 3, implementing $e^{i\Delta}$ by Hamiltonian simulation and phase estimating with input state $u_{k}$ is enough to estimate the Betti number $\beta_{k}$. This significantly reduces the complexity of computing the dual matrix in Eq. (10), where a complex norm matrix needs to be computed.
Theorem 3.
Using quantum phase estimation algorithm measuring the eigenvalues of $\Delta$ with the input state $u_{k}$, the Betti number $\beta_{k}$ of the GLMY homology can be estimated with the following procedure. The phase estimate procedure is repeated $M$ times and and the results of $M$ eigenvalues $\{\lambda_{i}\}_{i=1}^{M}$ are obtained. The relative frequency of the eigenvalue zero gives an estimate of the normalized Betti number $c_{k}$,
$c_{k}=\frac{\beta_{k}}{\textup{dim}\Gamma_{k}}=\frac{\textup{dim}\mathbf{H_{k}}}{\textup{dim}\Gamma_{k}}\approx\frac{|\{i,:\lambda_{i}=0\}|}{M},$ (15)
and the Betti number thus can be computed as
$\beta_{k}=c_{k}|\Gamma_{k}|.$ (16)
The proof of Theorem 3 is given in Appendix C.
III.1 Encoding of the Paths
Here we introduce an encoding protocol for encoding paths on directed graphs into quantum states. Suppose $p_{k}=v_{0}\cdots v_{k}$ where $v_{i}\in\{1,\cdots,n\}$ is a regular allowed path of length $k$, and $k\leq d$. Let $l(d)=\lceil\log(d+1)\rceil$, which is the smallest number of bits to represent the order information of vertex of the above paths. Then we will encode the path $p_{k}$ as an $(n\cdot l(d))$-qubit state, i.e.,
$|p_{k}\rangle\equiv|v_{0}\cdots v_{k}\rangle=\otimes_{i=1}^{n}|b^{i}_{1}\cdots b^{i}_{n(d)}\rangle,$ (17)
where if $i=v^{\alpha}$ with $\alpha\in\{0,\cdots,k\}$,
[…]
The input state $u_{k}$ (13) is prepared via two steps. In the first step, we will prepare the state
$|P_{k}\rangle=\frac{1}{\sqrt{|\Gamma_{k}|}}\sum_{p_{k}\in\Gamma_{k}}|p_{k}\rangle.$ (19)
In the second step, the mixed state $u_{k}=\frac{1}{|\Gamma_{k}|}\sum_{p_{k}\in\Gamma_{k}}|p_{k}\rangle\langle p_{k}|$ is constructed by using $n$-qubit ancillary qubits, applying a CNOT gate to each qubit in $|P_{k}\rangle$, and tracing out the ancillary qubits [17].
We introduce two methods to prepare $|P_{k}\rangle$. The first method is based on the Grover search algorithm, which is used in the previous LGZ algorithm [2]. This method begins from the initial state $|Q_{k}\rangle=\frac{1}{\sqrt{|\Lambda_{k}|}}\sum_{p_{k}\in\Lambda_{k}}|p_{k}\rangle.$ And the membership function is
$f_{k}(p_{k})=\begin{cases}1&\text{if }p_{k}\in\Gamma_{k},\\ 0&\text{if }p_{k}\notin\Gamma_{k}.\end{cases}$ (20)
The target state $|P_{k}\rangle$ can be arrived in $\mathcal{O}(\zeta_{k}^{-1/2})$ steps, where $\zeta_{k}=|\Gamma_{k}|/|\Lambda_{k}|$ is the fraction of the occupied $k$-paths in $\Gamma_{k}$ of all possible $k$-paths.
In the case where we can efficiently sample $k$-paths, we propose that we can use the the amplitude encoding [34, 35, 36, 37] to prepare the input state $|P_{k}\rangle$. The amplitude encoding is previously used for data embedding in QML [38]. The amplitude encoding is to prepare the state
$|\psi\rangle=\frac{1}{N}\sum_{i=0}^{N-1}c_{i}|i\rangle,$ (21)
where $N=\sum_{i}|c_{i}|^{2}$ is the normalization factor. In the case of preparing $|P_{k}\rangle$, the coefficients $c_{i}=1$ for all $i$, where $|i\rangle\in\Gamma_{k}$, and $\Gamma_{k}$ is the Fock space of molecular-system-like states with a conserved particle number. The scaling behavior for preparing $k$-path states is $n^{k}2^{k}/k!$, which is polynomially with $n$.
III.2 Encoding of Boundary Operator
Here we will first define the boundary map in our algorithm. The total boundary operator $\hat{\partial}=\sum_{k=0}^{d}\hat{\partial}_{k}$ for all paths in the digraph $G$ with
$\hat{\partial}_{k}$ $=\sum_{p_{k}\in\Lambda_{k}}|\partial p_{k}\rangle\langle p_{k}|$
$=\sum_{v_{0}\cdots v_{k}}\sum_{i=0}^{k}(-1)^{i}|v_{0}\cdots\hat{v}_{i}\cdots v_{k}\rangle\langle v_{0}\cdots v_{k}|$
$=\sum_{v_{0}\cdots v_{k}}\sum_{i=0}^{k}(-1)^{i}\otimes_{i=1}^{n}|\bar{a}^{i}\bar{b}_{1}^{i}\cdots\bar{b}_{n(d)}^{i}\rangle\langle a^{i}b_{1}^{i}\cdots b_{n(d)}^{i}|,$ (22)
where $b^{i}_{m}$ are specified by Eq.(18),
[…]
The total Dirac operator $B$ obtained from the total boundary $\partial$,
$B=\partial+\partial^{\dagger}.$ (24)
The matrix $B$ is $n[\log(d+1)]$-sparse, and the Hamiltonian simulation protocol can implement $e^{iB}$ using $\mathcal{O}(n^{3})$ gates. Denote the largest and the smallest eigenvalue of $B$ as $\lambda_{\text{max}}$ and $\lambda_{\text{min}}$, the Dirac operator needs to be rescaled by $\lambda_{\text{max}}^{-1}$, to avoid multiples of $2\pi$ in $e^{iB}$. And the phase estimate algorithm needs to be executed $\lambda^{-1}=\lambda_{\text{min}}/\lambda_{\text{{max}}}$ times. And the total complexity of phase estimate procedure is $\mathcal{O}(\lambda n^{3})$ [2, 16, 17].
Since $\Delta=B^{2}$, from the Theorem 2, the Hodge Laplacian on $\Gamma$ is
$\Delta^{\Gamma}=PBPBP,$ (25)
with $P=\sum_{k}P_{k}$ is the projector of the Hilbert space of all the $k$-paths states in $\Gamma_{k}$, and $P_{k}$ is the projector operator on the Hilbert space of $u_{k}$.
The final step is the phase estimation and measurements. The procedure is as the Theorem 3.
III.3 Complexity Analysis
In the problem of solving the Betti number of simplicial complex is defined as: given a clique complex defined by its vertices and edges and an integer $k\geq 0$ as input, output the $k$th Betti number $\beta_{k}$ of the clique complex. This problem is proved to be NP-hard [17]. The mainly limiting factor for the complexity is the Grover search of the input state of the $k$-path, which is the same as the previous LGZ algorithm. The Grover search requires time complexity of $\Omega\big{(}\sqrt{\tbinom{n}{k+1}/\dim\Gamma_{k}|}\big{)}$, and the lower bound of the phase estimation algorithm to compute the Betti number is $\Omega(\sqrt{\dim\Gamma_{k}/\beta_{k}})$. The total runtime is lower bounded by $\Omega\big{(}\sqrt{\tbinom{n}{k+1}/\beta_{k}}\big{)}$.
However, in the case where we can efficiently sample $k$-paths, by using the amplitude encoding to prepare the input states, the quantum algorithm realizes an exponential speedup. The lower bound of the for preparing the state $\rho_{k}$ is $n^{k}2^{k}/k!$, and the lower bound of the phase estimation algorithm to compute the Betti number is $\Omega(\dim\Gamma_{k}/\beta_{k})$. Combining with these two lower bound, the total lower bound of the quantum algorithm for GLMY homology is $\Omega\big{(}n^{k}2^{k}\dim\Gamma_{k}/k!\beta_{k}\big{)}$, which is polynomial in $n$.
IV Conclusion and Outlook
In this paper we propose the quantum algorithm for GLMY homology. We design a quantum state encoding protocol based on paths in the GLMY homology. And we design a efficient encoding protocol of the boundary map on quantum circuit. We prove that the null space of embedded homology equals to the null space of the the regular paths space projected on the embedded homology space. Finally, the quantum algorithm for GLMY homology gives an exponential quantum advantage in the case of the input states can be sampled efficiently, and it gives at least a quadratic speedup in other general cases.
Our encoding protocol can represent the directed edges on the quantum circuit, which generalizes the encoding methods from the previous simplicial complex. Our encoding protocol provides a framework for the further researches of the quantum algorithm on the digraph. There may have a further improvement of the method for finding the superposition of all $k$-th states. And a more efficient framework for the quantum algorithm of the topological data analysis may be proposed.
Acknowledgements.
This work is supported by Natural Science Foundation of Shandong Province (Grant No. ZR2025QC1476), National Key Research and Development Program of China (Grants No. 2021YFA1402104 and No.2021YFA0718302) and National Natural Science Foundation of China (Grant No. 12075310).
Appendix A Proof the Theorem 1
Proof.
(The Proof of Theorem 1) This theorem follows from a direct compute. Let us assume the dimensions of $\Lambda_{k}$, $\mathcal{A}_{k}$ and $\Gamma_{k}$ as $\lambda_{k}$, $\alpha_{k}$ and $\gamma_{k}$ respectively. A selection for the basis of $\Lambda_{k}$ is the tuple of all elementary regular $k$-paths which we denote it as $(R_{k})$. Clearly this is a orthogonal normal basis under the inner product $\langle-,-,\rangle$. Let $A_{k}$ be the set of all allowed elementary $k$-paths, the tuple $(A_{k},L_{k})$ given by adding the elements $L_{k}$ of $\partial^{\Gamma}_{k+1}A_{k+1}$ that are not belongs to $\mathcal{A}_{k}$ is the basis of $\Gamma_{k}$. Clearly $(A_{k},L_{k})$ is not orthogonal however it satisfies the properties
  • for any $x,y\in A_{k}$, $\langle x,y\rangle_{\Gamma}=1$ if and only if $x=y$ and 0 otherwise,
  • for $x\in A_{k}$ and $y\in L_{k}$, $\langle x,y\rangle_{\Gamma}=0$.
Let us assume the total boundary $\partial_{k}$ is given by the matrix $(D_{k})_{\lambda_{k-1}\times\lambda_{k}}$ and the embeding $\Gamma_{k}\hookrightarrow\Lambda_{k}$ is given by the matrix $(E_{k})_{\lambda_{k}\times\gamma_{k}}$ such that $\operatorname{rank}E_{k}=\gamma_{k}$. If the matrix of $\partial_{k}^{\Gamma}$ is assumed as $D_{k}^{\Gamma}$, we have
$\partial_{k}^{\Gamma}(A_{k},L_{k})$ $=$ $(A_{k-1},L_{k-1})D_{k}^{\Gamma},$ (26)
$\partial_{k}^{\Gamma}(A_{k},L_{k})$ $=$ $(R_{k-1})D_{k}E_{k},$
$(A_{k-1},L_{k-1})$ $=$ $(R_{k-1})E_{k-1}.$
Hence $D_{k}^{\Gamma}=(E_{k-1}^{\dagger}E_{k-1})^{-1}E_{k-1}^{\dagger}D_{k}E_{k}$.
The norm matrix of $\langle-,-\rangle_{\Lambda_{k}}$ is $\operatorname{Id}_{\lambda_{k}\times\lambda_{k}}$. This implies that the dual of $\partial_{k}^{*}$ is $D_{K}^{*}=D_{k}^{\dagger}$ where $D_{k}^{\dagger}$ is the transport of $D_{k}$. The norm matrix of $\langle-,-\rangle_{\Gamma_{k}}$ is $N_{k}=(E_{k}^{\dagger}E_{k})$. Therefore the dual $(\partial_{k}^{\Gamma})^{*}$ of $\partial_{k}^{\Gamma}$ under $\langle-,-\rangle_{\Gamma_{k}}$ is given as
$D_{k}^{\Gamma,*}=(E_{k}^{\dagger}E_{k})^{-1}E_{k}^{\dagger}D_{k}^{\dagger}E_{k-1}$ (27)
On the other hand, the projection $p_{k}:\Lambda_{k}\to\Lambda_{k}$ onto the $\Gamma_{k}$ is given by the matrix $P_{k}=E_{k}(E_{k}^{\dagger}E_{k})^{-1}E_{k}^{\dagger}$. The map $p_{k}\circ\partial_{k}^{*}|_{\Gamma_{k}}$ is given as
$\partial_{k}^{*}\circ e_{k-1}=p_{k}\circ\partial_{k}^{*}|_{\Gamma_{k-1}}=E_{k}(E_{k}^{\dagger}E_{k})^{-1}E_{k}^{\dagger}D_{k}^{\dagger}E_{k-1}=e_{k}\circ(\partial_{k}^{\Gamma})^{*}$ (28)
This finishes the proof. ∎
Summarize the Equation 4 and Theorem 1 we have the following commutative diagram
[…]
Appendix B Proof the Theorem 2
Proof.
(The Proof of Theorem 2) It is clear from the definition that $p_{k}\circ\Delta_{k}|_{\Gamma_{k}}=p_{k}\circ\Delta_{k}\circ e_{k}$. Then from the Diagram 29, we have the following equations
$p_{k}\circ\Delta_{k}\circ e_{k}$ $=$ $p_{k}\circ\partial_{k}^{*}\circ\partial_{k}\circ e_{k}+p_{k}\circ\partial_{k+1}\circ\partial_{k+1}^{*}\circ e_{k}$ (30)
$=$ $p_{k}\circ\partial_{k}^{*}\circ e_{k-1}\circ\partial_{k}^{\Gamma}+p_{k}\circ\partial_{k+1}\circ e_{k+1}\circ(\partial_{k+1}^{\Gamma})^{*}$
$=$ $p_{k}\circ\partial_{k}^{*}\circ e_{k-1}\circ\partial_{k}^{\Gamma}+p_{k}\circ e_{k}\circ\partial_{k+1}^{\Gamma}\circ(\partial_{k+1}^{\Gamma})^{*}$
$=$ $(\partial_{k}^{\Gamma})^{*}\partial_{k}^{\Gamma}+\partial_{k+1}^{\Gamma}(\partial_{k+1}^{\Gamma})^{*}=\Delta_{k}^{\Gamma}$
Hence the operator $\Delta_{k}^{\Gamma}$ has the same nullspace with the operator $p_{k}\circ\Delta_{k}|_{\Gamma_{k}}$. ∎
Appendix C Proof the Theorem 3
Proof.
(The Proof of Theorem 3) We focus on computing the $k$th Betti number $\beta_{k}$. The input state is
$u_{k}=\frac{1}{\dim\Gamma_{k}}\sum_{p_{k}\in\Gamma_{k}}|p_{k}\rangle\langle p_{k}|.$ (31)
The Hodge Laplacian $\Delta^{\Gamma}$ on $\Gamma$ is a direct sum of $k$-th terms
$\Delta^{\Gamma}=\mathop{\bigoplus}\limits_{k=0}^{n-1}\Delta^{\Gamma}_{k}.$ (32)
Let us focus on the $k$th sector because the state $u_{k}$ is only in the space of $\Delta^{\Gamma}_{k}$. The matrix $\Delta^{\Gamma}_{k}$ can be decomposed into two parts,
$\Delta^{\Gamma}_{k}=\sum_{i}a^{\Gamma}_{i}|k^{\Gamma}_{i}\rangle\langle k^{\Gamma}_{i}|+0\sum_{p^{0}_{k}\in\Gamma_{k}}|p^{0}_{k}\rangle\langle p^{0}_{k}|,$ (33)
where $|k_{i}^{\Gamma}\rangle$ is the eigenvector of $\Delta^{\Gamma}_{k}$ with non-zero eigenvalue $a_{i}$, and $|p^{0}_{k}\rangle$ is the quantum state of $k$-path simplex in the set of $\Gamma_{k}$ with zero eigenvalue. The $\Delta$ also has a direct sum structure, and can be decomposed into two parts, $\Delta=\bigoplus_{k=0}^{n-1}\Delta_{k}$. We only focus on the $k$th sector,
$\Delta_{k}=\sum_{i}a_{i}|k_{i}\rangle\langle k_{i}|+0\sum_{q^{0}_{k}\in\Lambda_{k}}|q^{0}_{k}\rangle\langle q^{0}_{k}|,$ (34)
where $|q^{0}_{k}\rangle$ is the quantum state of $k$-path simplex in the set of $\Lambda_{k}$ with zero eigenvalue.
We assert that taking the input state $u_{k}$ and simulating the Hamiltonian $\Delta$ in the phase estimate algorithm is equivalent to simulating the Hamiltonian $P^{\Gamma}_{k}\Delta P^{\Gamma}_{k}$. Recall that in the phase estimate algorithm, the probability of getting an eigenvalue $a_{i}$ of $\Delta$ is equal to the overlap of the $u_{k}$ and the corresponding eigenvectors $|k_{i}\rangle$ of the Hamiltonian $\Delta_{k}$ by implement the evolution $e^{i\Delta_{k}}$ and measuring the eigenvalues with the input state $u_{k}$.
$p(a_{i})$ $=\langle k_{i}|u_{k}|k_{i}\rangle$
$=\langle k_{i}|P^{\Gamma}_{i}u_{k}P^{\Gamma}_{i}|k_{i}\rangle.$ (35)
The second equality is obtained by the $P_{k}$ is the projector of $u_{k}$, i.e., $P^{\Gamma}_{k}u_{k}P^{\Gamma}_{k}=u_{k}$. Note that
$P^{\Gamma}_{k}\Delta P^{\Gamma}_{k}=\sum_{i}a_{i}P^{\Gamma}_{k}|k_{i}\rangle\langle k_{i}|P^{\Gamma}_{k},$ (36)
where $P^{\Gamma}_{k}$ is the projection operator of the space of $\Gamma_{k}$. Thus taking the input state $u_{k}$ and simulating the Hamiltonian $\Delta$ in the phase estimate algorithm is equivalent to simulating the Hamiltonian $P^{\Gamma}_{k}\Delta P^{\Gamma}_{k}$. From Theorem 2, $\beta_{k}=\text{dim}~\text{Ker}\Delta^{\Gamma}=\text{dim}~\text{Ker}P^{\Gamma}_{k}\Delta P^{\Gamma}_{k}$. So the Betti number $\beta_{k}$ can be obtained by simulating $e^{i\Delta}$ and phase estimating with the input state $u_{k}$.
For the measurements of zero eigenvalues,
$\dim\Gamma_{k}\sum_{i}p(0_{i})$ $=\dim\Gamma_{k}\bigg{[}\sum_{p^{0}_{k}\in\Gamma_{k}}\langle p^{0}_{k}|\bigg{(}\frac{1}{\dim\Gamma_{k}}\sum_{p_{k}\in\Gamma_{k}}|p_{k}\rangle\langle p_{k}|\bigg{)}|p^{0}_{k}\rangle\bigg{]}$
$=\#\{0\text{-eigenvalue}\}$
$=\text{dim}~\text{Ker}\Delta_{k}^{\Gamma}.$ (37)
The $k$th Betti number $\beta_{k}$ can be obtained from the relative frequency of the measurement of a zero eigenvalue. The normalized Betti number $\beta_{k}$ is,
$c_{k}=\frac{\beta_{k}}{\textup{dim}\Gamma_{k}}=\frac{\textup{dim}\mathbf{H_{k}}}{\textup{dim}\Gamma_{k}}\approx\frac{|\{i,:\lambda_{i}=0\}|}{M}.$ (38)
And the Betti number thus can be computed as
$\beta_{k}=c_{k}|\Gamma_{k}|.$ (39)
Appendix D Examples of Classical Computation of GLMY Homology
In this section we will give a direct computation for the terms, boundary maps in the sequence of Eq. (40) and the Betti numbers.
$\cdots\stackrel{{\scriptstyle\partial^{\Gamma}_{5}}}{{\longrightarrow}}\Gamma_{4}\stackrel{{\scriptstyle\partial^{\Gamma}_{4}}}{{\longrightarrow}}\Gamma_{3}\stackrel{{\scriptstyle\partial^{\Gamma}_{3}}}{{\longrightarrow}}\Gamma_{2}\stackrel{{\scriptstyle\partial^{\Gamma}_{2}}}{{\longrightarrow}}\Gamma_{1}\stackrel{{\scriptstyle\partial^{\Gamma}_{1}}}{{\longrightarrow}}\Gamma_{0}\stackrel{{\scriptstyle\partial^{\Gamma}_{0}}}{{\longrightarrow}}0.$ (40)
Example 1.
Let $G$ be the digraph as the Fig. 1 (a).
Figure 1: Two examples of digraphs.
We will show all its Betti numbers are zero except $\beta_{0}$. It is clear the group of allowed paths are as following
$\mathcal{A}_{0}$ $=$ $\operatorname{span}\{$ $1,2,3,4\},$ (41)
[…]
And for $k\geq 4$, the group of allowed paths $\mathcal{A}_{k}=0$ all vanish. The $k$-th chain groups $\Gamma_{k}=\mathcal{A}_{k}=0$ and $k$-th boundary maps $\partial_{k}=\partial_{k}^{\Gamma}=0$. When $k=0$, the chain group $\Gamma_{0}$ is always equal to the group of all vertices, i,e $\Gamma_{0}=\mathcal{A}_{0}=\operatorname{span}\{1,2,3,4\}$. The norm matrix of $\langle-,-\rangle_{\Gamma_{0}}$ is $\operatorname{Id}_{4\times 4}$. And the 0-th boundary map $\partial_{0}=\partial^{\Gamma}_{0}=0$.
For $1\leq k\leq 3$, from the definition, the boundary operators are listed as following
[…]
Therefore, the third chain group $\Gamma_{3}=\mathcal{A}_{3}+\partial_{4}\mathcal{A}_{4}=\operatorname{span}\{1234\}$. The norm matrix of $\langle-,-\rangle_{\Gamma_{3}}$ is $N_{3}=1$. The second chain group $\Gamma_{2}=\mathcal{A}_{2}+\partial_{3}\mathcal{A}_{3}=\operatorname{span}\{123,124,134,234\}=\mathcal{A}_{2}$. The norm matrix of $\langle-,-\rangle_{\Gamma_{2}}$ is $N_{2}=\operatorname{Id}_{4\times 4}$. And the first chain group $\Gamma_{1}=\mathcal{A}_{1}+\partial_{2}\mathcal{A}_{2}=\operatorname{span}\{12,13,14,23,24,34\}=\mathcal{A}_{1}$ and the norm matrix of $\langle-,-\rangle_{\Gamma_{1}}$ is $N_{1}=\operatorname{Id}_{6\times 6}$
Then the boundary matrices $D_{k}^{\Gamma}$ of the boundary maps $\partial^{\Gamma}_{k}$ when $k=1,2,3$ are listed as following
[…]
$=$ $(1,2,3,4)D_{1}^{\Gamma}$
$=$ $(1,2,3,4)\begin{pmatrix}-1&-1&-1&0&0&0\\ 1&0&0&-1&-1&0\\ 0&1&0&1&0&-1\\ 0&0&1&0&1&1\\ \end{pmatrix}$
Since all the norm matrix are identity, computation shows that the Hodge-Laplacian operators are as the equation 44. It is clear that $\Delta^{\Gamma}_{3},\ \Delta^{\Gamma}_{2}$ and $\Delta^{\Gamma}_{1}$ are full rank while $\operatorname{rank}\Delta^{\Gamma}_{0}=3$. Hence the Betti numbers are $\beta_{0}=1$ and the others are 0.
$\Delta^{\Gamma}_{3}$ $=$ $(D_{3}^{\Gamma})^{\dagger}D_{3}^{\Gamma}$ $=$ $4\operatorname{Id}_{1\times 1}$ (44)
$\Delta^{\Gamma}_{2}$ $=$ $D_{3}^{\Gamma}(D_{3}^{\Gamma})^{\dagger}+(D_{2}^{\Gamma})^{\dagger}D_{2}^{\Gamma}$ $=$ $4\operatorname{Id}_{4\times 4}$
$\Delta^{\Gamma}_{1}$ $=$ $D_{2}^{\Gamma}(D_{2}^{\Gamma})^{\dagger}+(D_{1}^{\Gamma})^{\dagger}D_{1}^{\Gamma}$ $=$ $4\operatorname{Id}_{6\times 6}$
$\Delta^{\Gamma}_{0}$ $=$ $D_{1}^{\Gamma}(D_{1}^{\Gamma})^{\dagger}$ $=$ $\begin{pmatrix}3&-1&-1&-1\\ -1&3&-1&-1\\ -1&-1&3&-1\\ -1&-1&-1&3\\ \end{pmatrix}$
Example 2.
Let $G$ be the digraph as the Figure 1 (b).
It was proved in [1] that this digraph has the properties $\beta_{0}=\beta_{1}=1$ and $\beta_{k}=0$for any $k\geq 2$. We will check these properties by direct computations.
$\mathcal{A}_{0}$ $=$ $\operatorname{span}\{$ $0,1,2,3,4,5\},$ (45)
[…]
And for $k\geq 3$, the group of allowed paths $\mathcal{A}_{k}=0$ all vanish. The $k$-th chain groups $\Gamma_{k}=\mathcal{A}_{k}=0$ and $k$-th boundary maps $\partial_{k}=\partial_{k}^{\Gamma}=0$. When $k=0$, the chain group $\Gamma_{0}$ is always equal to the group of all vertices, i,e $\Gamma_{0}=\mathcal{A}_{0}=\operatorname{span}\{0,1,2,3,4,5\}$. The norm matrix of $\langle-,-\rangle_{\Gamma_{0}}$ is $\operatorname{Id}_{6\times 6}$. And the 0-th boundary map $\partial_{0}=\partial^{\Gamma}_{0}=0$.
When $k=1,2$, the boundary maps are given as
[…]
Hence the second chain group $\Gamma_{2}=\mathcal{A}_{2}+\partial_{3}\mathcal{A}_{3}=\operatorname{span}\{013,014,023,024\}$. The norm matrix of $\langle-,-\rangle_{\Gamma_{2}}$ is $N_{2}=\operatorname{Id}_{4\times 4}$. The first chain group $\Gamma_{1}=\mathcal{A}_{1}+\partial_{2}\mathcal{A}_{2}=\operatorname{span}\{01,02,13,14,23,24,53,54,03,04\}$, where $03,04$ are generated from $\partial_{2}\mathcal{A}_{2}$. The norm matrix of $\langle-,-\rangle_{\Gamma_{1}}$ is $N_{1}=\operatorname{Id}_{10\times 10}$.
The computation shows that the boundary matrices $D_{2}^{\Gamma}$ resp. $D_{1}^{\Gamma}$ of the boundary maps $\partial^{\Gamma}_{2}$ resp. $\partial^{\Gamma}_{1}$ are as following
[…]
$=$ $(0,1,2,3,4,5)D_{1}^{\Gamma}$
$=$ $(0,1,2,3,4,5)\left(\begin{array}[]{cccccccccc}-1&-1&0&0&0&0&0&0&-1&-1\\ 1&0&-1&-1&0&0&0&0&0&0\\ 0&1&0&0&-1&-1&0&0&0&0\\ 0&0&1&0&1&0&1&0&1&0\\ 0&0&0&1&0&1&0&1&0&1\\ 0&0&0&0&0&0&-1&-1&0&0\\ \end{array}\right)$
Finally the Hodge-Laplacian operators are as the equation 48. The Betti numbers are $\beta_{0}=\beta_{1}=1$ where $\ker\Delta^{\Gamma}_{1}$ is generated by (01+1423+24+35335403+04)(-01+14-23+24+3*53-3*54-03+04) and $\ker\Delta^{\Gamma}_{0}$ is generated by $(0+1+2+3+4+5)$. In other cases, the Betti numbers are all zero.
$\Delta^{\Gamma}_{2}$ $=$ $(D_{2}^{\Gamma})^{\dagger}D_{2}^{\Gamma}$ $=$ $\left(\begin{array}[]{cccc}3&1&1&0\\ 1&3&0&1\\ 1&0&3&1\\ 0&1&1&3\\ \end{array}\right)$ (48)
$\Delta^{\Gamma}_{1}$ $=$ $D_{2}^{\Gamma}(D_{2}^{\Gamma})^{\dagger}+(D_{1}^{\Gamma})^{\dagger}D_{1}^{\Gamma}$ $=$ $\left(\begin{array}[]{cccccccccc}4&1&0&0&0&0&0&0&0&0\\ 1&4&0&0&0&0&0&0&0&0\\ 0&0&3&1&1&0&1&0&0&0\\ 0&0&1&3&0&1&0&1&0&0\\ 0&0&1&0&3&1&1&0&0&0\\ 0&0&0&1&1&3&0&1&0&0\\ 0&0&1&0&1&0&2&1&1&0\\ 0&0&0&1&0&1&1&2&0&1\\ 0&0&0&0&0&0&1&0&4&1\\ 0&0&0&0&0&0&0&1&1&4\\ \end{array}\right)$
$\Delta^{\Gamma}_{0}$ $=$ $D_{1}^{\Gamma}(D_{1}^{\Gamma})^{\dagger}$ $=$ $\left(\begin{array}[]{cccccc}4&-1&-1&-1&-1&0\\ -1&3&0&-1&-1&0\\ -1&0&3&-1&-1&0\\ -1&-1&-1&4&0&-1\\ -1&-1&-1&0&4&-1\\ 0&0&0&-1&-1&2\\ \end{array}\right)$

### v2

@@
 And a property of the GLMY homology is proved for the theoretical guarantee of the quantum algorithm.
-The quantum algorithm for GLMY homology gives a quadratic speedup in general cases, and it gives an exponential quantum advantage in the case of the input data is given as a specification of paths.
+The quantum algorithm for GLMY homology provides a cubic speedup in general cases, and it can provide an exponential quantum advantage in the case of the input data is given as a specification of paths.
 I Introduction
@@
 By applying the topological tools, for example the simplicial homology theory and the knot theory etc., it is able to analyze complex high-dimensional data and identify key features of the data which may be difficult to be extracted with traditional methods.
-Quantum algorithm for TDA is first proposed by Lloyd, Garnerone, and Zanardi (LGZ) [2], and it gives significant advantage over the best classical algorithm.
+Quantum algorithm for TDA is first proposed by Lloyd, Garnerone, and Zanardi (LGZ) [2], and it provides significant advantage over the best classical algorithm.
 The foundations of the LGZ algorithm are linear-algebraic quantum machine learning (QML) algorithms [3, 4, 5, 6, 7, 8, 9, 10], which are one of the most rapidly growing fields in the research of quantum computing.
 Recent studies on quantum algorithms for TDA have achieved notable theoretical progress [11, 12, 13, 14, 15, 16, 17, 18].
-The complexity of the LGZ algorithm is analyzed in [17] and the conclusion is that computing Betti numbers is NP-hard and it achieves quadratic speedup for inputs as vertices and edges.
+The complexity of the LGZ algorithm is analyzed in [17] and the conclusion is that computing Betti numbers is NP-hard and it achieves cubic speedup for inputs as vertices and edges.
 It can achieve exponential advantage only when inputs are given as a simplex list.
@@
 Furthermore, a property of the GLMY homology is proved for the theoretical guarantee of the quantum algorithm thus the norm matrix used in classical algorithm can be ignored.
-Given the input data as a specification of paths, the amplitude encoding method can be used to efficiently prepare the input state.
-The quantum algorithm for GLMY homology gives a quadratic speedup in general cases, and it gives an exponential quantum advantage in the special case of the input data is given as a specification of paths that can be sampled efficiently.
+In the case where the input data as a specification of paths is given as the input, the amplitude encoding method can be used to efficiently prepare the input state.
+The quantum algorithm for GLMY homology provides a cubic speedup in the case where the input is given as vertices and edges, and it can provide an exponential quantum advantage in the special case where the input data is given as a specification of paths.
 II GLMY Homology: Definitions and Computations
@@
 The target state $|P_{k}\rangle$ can be arrived in $\mathcal{O}(\zeta_{k}^{-1/2})$ steps, where $\zeta_{k}=|\Gamma_{k}|/|\Lambda_{k}|$ is the fraction of the occupied $k$-paths in $\Gamma_{k}$ of all possible $k$-paths.
-In the case where we can efficiently sample $k$-paths, we propose that we can use the the amplitude encoding [34, 35, 36, 37] to prepare the input state $|P_{k}\rangle$.
+In the case where in the special case where the input data is given as a specification of paths, we propose that we can use the the amplitude encoding [34, 35, 36, 37, 38] to prepare the input state $|P_{k}\rangle$.
 The amplitude encoding is previously used for data embedding in QML [39].
@@
 In the case of preparing $|P_{k}\rangle$, the coefficients $c_{i}=1$ for all $i$, where $|i\rangle\in\Gamma_{k}$, and $\Gamma_{k}$ is the Fock space of molecular-system-like states with a conserved particle number.
-The scaling behavior for preparing $k$-path states is $n^{k}2^{k}/k!$, which is polynomially with $n$.
+The input state $|P_{k}\rangle$ can be prepared in $|\Gamma_{k}|^{2}(n\log n)^{2}$ steps [38], where $n\log n$ is the qubits number of the state $|P_{k}\rangle$.
 III.2 Encoding of Boundary Operator
@@
 The mainly limiting factor for the complexity is the Grover search of the input state of the $k$-path, which is the same as the previous LGZ algorithm.
-The Grover search requires time complexity of $\Omega\big{(}\sqrt{\tbinom{n}{k+1}/\dim\Gamma_{k}|}\big{)}$, and the lower bound of the phase estimation algorithm to compute the Betti number is $\Omega(\sqrt{\dim\Gamma_{k}/\beta_{k}})$.
+The Grover search requires time complexity of $\Omega\big(\sqrt{\tbinom{n}{k+1}/|\Gamma_{k}}|\big)$, and the lower bound of the phase estimation algorithm to compute the Betti number is $\Omega\big(\sqrt{|\Gamma_{k}|/\beta_{k}}\big)$.
 The total runtime is lower bounded by $\Omega\big(\sqrt{\tbinom{n}{k+1}/\beta_{k}}\big)$.
-However, in the case where we can efficiently sample $k$-paths, by using the amplitude encoding to prepare the input states, the quantum algorithm realizes an exponential speedup.
-The lower bound of the for preparing the state $\rho_{k}$ is $n^{k}2^{k}/k!$, and the lower bound of the phase estimation algorithm to compute the Betti number is $\Omega(\dim\Gamma_{k}/\beta_{k})$.
-Combining with these two lower bound, the total lower bound of the quantum algorithm for GLMY homology is $\Omega\big{(}n^{k}2^{k}\dim\Gamma_{k}/k!\beta_{k}\big{)}$, which is polynomial in $n$.
+However, in the case where the input is given by the list of all $k$-path, by using the amplitude encoding to prepare the input states, the quantum algorithm can realize an exponential speedup.
+The complexity for preparing the state $\rho_{k}$ is $|\Gamma_{k}|^{2}(n\log n)^{2}$.
+In the case where $\Gamma_{k}$ is sparse, $|\Gamma_{k}|^{2}(n\log n)^{2}$ is scaling polynomially.
+Combining with the lower bound of the phase estimation, the total lower bound of the quantum algorithm for GLMY homology is $\Omega\big((n\log n)^{2}\sqrt{|\Gamma_{k}|^{5}/\beta_{k}}\big)$, which achieves an exponential speedup.
 IV Conclusion and Outlook
@@
 We design a quantum state encoding protocol based on paths in the GLMY homology.
-And we design a efficient encoding protocol of the boundary map on quantum circuit.
 We prove that the null space of embedded homology equals to the null space of the the regular paths space projected on the embedded homology space.
-Finally, the quantum algorithm for GLMY homology gives an exponential quantum advantage in the case of the input states can be sampled efficiently, and it gives at least a quadratic speedup in other general cases.
+Finally, the quantum algorithm for GLMY homology can provide an exponential quantum advantage in the case of the input states can be sampled efficiently, and it provides at least a cubic speedup in the case where the input is given as vertices and edges.
 Our encoding protocol can represent the directed edges on the quantum circuit, which generalizes the encoding methods from the previous simplicial complex.
@@
 Acknowledgements.
-This work is supported by Natural Science Foundation of Shandong Province (Grant No.
-ZR2025QC1476), National Key Research and Development Program of China (Grants No.
-2021YFA1402104 and No.2021YFA0718302) and National Natural Science Foundation of China (Grant No.
-12075310).
+This work is supported by Science Challenge Project (Grant No.
+TZ2025017), National Key Research and Development Program of China (Grants No.
+2021YFA1402104 and No.2021YFA0718302), National Natural Science Foundation of China (Grant No.
+12075310) and Natural Science Foundation of Shandong Province (Grant No.
+ZR2025QC1476).
 Appendix A Proof the Theorem 1